OpenStudy (yacoub1993):

Simplify the expression.3/a^-9b

OpenStudy (yacoub1993):

\[\frac{ 3 }{ a ^{-9}b }\]

OpenStudy (yacoub1993):

@campbell_st is this correct answer \[\frac{ 3a }{ b }\]

OpenStudy (saifoo.khan):

You forgot to add 9.

OpenStudy (saifoo.khan):

a^9

OpenStudy (saifoo.khan):

Now you forgot to add a 'b'. :P

OpenStudy (saifoo.khan):

Again. :P

OpenStudy (yacoub1993):

\[\frac{ 3a ^{9} }{ b }\]

OpenStudy (anonymous):

3/a-9b = 3(1-3ab)/a

OpenStudy (yacoub1993):

@evelynn92 how?

OpenStudy (yacoub1993):

the answer i posted is correct...

OpenStudy (yacoub1993):

@saifoo.khan

OpenStudy (anonymous):

3a^9/b

OpenStudy (anonymous):

Ur correct

OpenStudy (saifoo.khan):

Yes. You're right.